1.A A mode choice study resulted in the following utility equation UASC Where -0.025AETm -0.032WT-0.015LHT -0.002T ASC Altemative Specific Constant, 0 for Auto and -0.1 for bus AET-Access plus Egress Time for mode m (minutes) WT- Wait Time for mode m (m inutes) LHT, Line Haul Time for mode m (minutes) TC Out of pocket costs for mode m, for the bus this is the bus fare (cents) The trip forecast for a particular target-year was Tu -5000 person-trips per day. During the target year trip- makers will have a choice between the private automobile (A) and a local bus system (B). The target-year service attributes of the two competing modes have been estimated to be Attribute Automobile Local bus WT 0 15 LHT 20 40 AET TC 100 50 10 a) Apply the logit model to estimate the target-year market share of the two modes and the resulting fare- box revenue of the bus system b) What are the values of (i) WT (ii) LHT (İİİ) AET for travelers? (in dollars per hour)

Urban Planning For Engineers Civil Engineering Mode Choice Utility Equation Logit Model
Answer #1

Solution:

For the private automobile A,

Um = ASC - 0.025 AET - 0.032 WT - 0.015 LHT - 0.002 TC

for auto, ASC = 0,

putting the given values of the table we get,

Uauto = 0 -0.025*5 - 0.032*0 - 0.015*20 - 0.002*100 = - 0.625

Ubus = -0.1 - 0.025*10 - 0.032*15 - 0.015*40 - 0.002*50 = -1.53

Probability = P

Pauto = e^(Uauto) / [e^Uauto + e^Ubus ]

= e^-0.625 / (e^-0.625 + e^-1.53 )

= 0.71198

proportion of trips by auto = 5000 * 0.71198 = 3560

Pbus =e^(Ubus) / [e^Uauto + e^Ubus ]

= e^-1.53 / (e^-0.625 + e^-1.53 )

= 0.28802

proportion of trips by bus = 1440

fare box revenue for bus = 1440 * 50 = 72000 cents = $ 720